### v2

# A False Discovery Rate Control Method Using a Fully Connected Hidden Markov Random Field for Neuroimaging Data

Abstract
False discovery rate (FDR) control methods are essential for voxel-wise multiple testing in neuroimaging data analysis, where hundreds of thousands or even millions of tests are conducted to detect brain regions associated with disease-related changes. Classical FDR control methods (e.g., BH, q-value, and LocalFDR) assume independence among tests and often lead to high false non-discovery rates (FNR). Although various spatial FDR control methods have been developed to improve power, they still fall short of jointly addressing three major challenges in neuroimaging applications: capturing complex spatial dependencies, maintaining low variability in both false discovery proportion (FDP) and false non-discovery proportion (FNP) across replications, and achieving computational scalability for high-resolution data. To address these challenges, we propose fcHMRF-LIS, a powerful, stable, and scalable spatial FDR control method for voxel-wise multiple testing. It integrates the local index of significance (LIS)-based testing procedure with a novel fully connected hidden Markov random field (fcHMRF) designed to model complex spatial structures using a parsimonious parameterization. We develop an efficient expectation-maximization algorithm incorporating mean-field approximation, the Conditional Random Fields as Recurrent Neural Networks (CRF-RNN) technique, and permutohedral lattice filtering, reducing the time complexity from quadratic to linear in the number of tests. Extensive simulations demonstrate that fcHMRF-LIS achieves accurate FDR control, lower FNR, reduced variability in FDP and FNP, and a higher number of true positives compared to existing methods. Applied to an FDG-PET dataset from the Alzheimer’s Disease Neuroimaging Initiative, fcHMRF-LIS identifies neurobiologically relevant brain regions and offers notable advantages in computational efficiency.
keywords:
Multiple testing , Neuroimaging , False discovery rate , False non-discovery rate , Fully connected graph , CRF-RNN
journal: arXiv
[…]
[label1]organization=Department of Biostatistics, School of Global Public Health, New York University,city=New York,state=NY,country=USA
[…]
[label2]organization=Department of Population and Public Health Sciences, University of Southern California,city=Los Angeles,state=CA,country=USA
[…]
[label3]organization=Brain Health Imaging Institute, Department of Radiology, Weill Cornell Medicine,city=New York,state=NY,country=USA
1 Introduction
Voxel-wise multiple testing is indispensable in neuroimaging data analyses, which often involve hundreds of thousands or even millions of voxel-level comparisons to identify brain regions that exhibit statistically significant differences between population groups [1, 2, 3, 4]. For example, in Alzheimer’s disease (AD) research, Fluorine-18 fluorodeoxyglucose positron emission tomography (FDG-PET) is widely used to measure brain glucose metabolism, serving as a neurodegeneration biomarker to support clinical diagnosis and track disease progression [5]. To uncover disease-related functional abnormalities, FDG-PET studies commonly conduct voxel-wise multiple testing across groups at different AD-related stages to localize brain regions of altered metabolic activity [6, 7, 8].
In such high-dimensional settings, multiple testing methods typically aim to control the false discovery rate (FDR), a measure of type I error, defined as the mean of the false discovery proportion (FDP). Compared with family-wise error rate control, FDR control generally offers higher statistical power in large-scale multiple testing problems [9]. However, when applied to spatial data, classical FDR control methods, such as BH [9], q-value [10], and LocalFDR [11], ignore spatial dependence by treating tests as independent, which results in a substantial loss of statistical power [12]. To mitigate this issue, various spatial FDR control methods have been developed. These include random field-based approaches [13, 12, 14, 15], local smoothing-based procedures [16, 17, 18, 19], and deep learning-based methods [20, 21], although not all were originally designed for neuroimaging data. Despite these advances, current spatial FDR control methods still fall short of jointly addressing the following three major challenges in neuroimaging applications.
First, existing spatial FDR control methods often fail to adequately capture the complex spatial dependencies inherent in neuroimaging data, such as spatial heterogeneity [22], distance-related dependence [23], and long-range interactions [24]. Examples include methods that assume homogeneous hidden Markov random fields (HMRFs) [15] or homogeneous Gaussian random fields [14], rely on (hidden) Markov random fields with short-range neighborhoods [13, 12, 15], or employ local smoothing approaches [16, 17, 18, 19]. This raises concerns about the validity of such methods in controlling the FDR or their optimality in minimizing the false non-discovery rate (FNR) [25], a measure of type II error, defined as the mean of false non-discovery proportion (FNP).
Figure 1: Comparison of nine FDR control methods under the simulation setting in Section 3.2 with $(\mu_{1},\sigma_{1}^{2})=(-2,1)$, approximately 30% signal proportion, and a nominal FDR level of 0.05.
Second, while state-of-the-art methods focus on controlling the FDR (the mean of FDP) and some also aim to minimize the FNR (the mean of FNP), they often exhibit high variability in FDP or FNP, leading to instability in results across replications. Figure 1 illustrates the variability in FDP, FNP, and the number of true positives (TP) detected across 50 simulation replications, comparing nine FDR control methods: three non-spatial methods, BH [9], q-value [10], and LocalFDR [11]; nnHMRF-LIS [12], which uses a nearest-neighbor HMRF (nnHMRF); two local smoothing-based methods, SmoothFDR [17] and LAWS [18]; two deep learning-based methods, NeuralFDR [20] and DeepFDR [21]; and our proposed method. Notably, although the two most recent methods LAWS and DeepFDR empirically control the FDR at the nominal level of 0.05, their standard deviations (SDs) of FDP are large, exceeding three times that of BH, and they also show unexpectedly high SDs of FNP and TP. This instability is particularly concerning in neuroimaging, where multiple testing is inherently exploratory and the truth status of each hypothesis is unobservable due to the unsupervised nature of the task. If spatial FDR control methods yield inconsistent results across experiments, their reliability and practical utility in neuroimaging applications are seriously undermined.
Third, most spatial FDR control methods (e.g., [12, 14, 20, 17, 18, 19]) lack computational scalability, rendering them computationally expensive for analyzing large-scale, high-resolution neuroimaging datasets. For example, in our real data analysis (Section 4) using FDG-PET data with 439,758 brain voxels of interest, LAWS [18] required an average of 149.5 hours on 20 CPU cores, while the deep learning-based method NeuralFDR [20] still needed 25.7 hours even when run on a GPU.
To address the three challenges, we propose fcHMRF-LIS, a novel spatial FDR control method for voxel-wise multiple testing in neuroimaging data, which aims to control the FDR, minimize the FNR, and maintain low variability in both FDP and FNP. The fcHMRF-LIS method builds on the local index of significance (LIS)-based testing procedure [26], implemented under our proposed fully connected HMRF (fcHMRF) model.
The LIS-based testing procedure [26] is a widely used framework for FDR control when tests are dependent. The LIS serves as a substitute for the traditional p-value. For each test, the LIS is defined as the conditional probability that the null hypothesis is true given the test statistics from all tests. Under mild conditions, the LIS-based testing procedure asymptotically minimizes the FNR while controlling the FDR at a prespecified level [26, 27, 12]. The key challenge in applying this procedure lies in accurately modeling the dependence among tests in order to estimate the unknown LIS values in practice. Previous works have explored this challenge in various data settings. For one-dimensional data, such as influenza-like illness surveillance time series or single nucleotide polymorphisms (SNPs) on chromosomes in genome-wide association studies (GWAS), hidden Markov chain models have been used to estimate LIS values [26, 28, 29]. In GWAS, other studies have employed HMRFs to model spatial dependence among SNPs using prior linkage disequilibrium information to define graph structures [30, 31]. In neuroimaging, nnHMRF-LIS [12] and its variant [15] use HMRFs with short-range neighborhoods, while DeepFDR [21] adopts a deep neural network called W-Net to learn complex spatial dependence.
However, as shown in Figure 1, neither nnHMRF-LIS nor DeepFDR perform well in our simulation study. The failure of nnHMRF-LIS to control the FDR stems from its reliance on a short-range neighborhood structure, which is inadequate for capturing the complex spatial dependencies in neuroimaging data, the first major challenge outlined earlier. DeepFDR, on the other hand, suffers from substantial variability across replications, likely due to overfitting. Although it adopts a sophisticated W-Net architecture with millions of parameters, in the multiple testing setting, DeepFDR uses only a single image of test statistics and corresponding p-values as input, providing limited information for training. In contrast, nnHMRF-LIS exhibits low variability, probably owing to its parsimonious model with only a few parameters. This observation motivates our extension of nnHMRF to an fcHMRF model that retains a small number of parameters for low variability while effectively capturing complex neuroimaging spatial dependencies.
Our proposed fcHMRF is inspired by the fully connected conditional random field (CRF) with appearance and smoothness kernels, a powerful model widely used in image segmentation [32, 33, 34]. In a fully connected CRF, each voxel is represented as a node in an undirected graph, and all pairs of nodes are connected by edges, which allows the model to capture long-range interactions. The appearance and smoothness kernels further enable the model to account for spatial heterogeneity and distance-related dependence. We innovatively formulate our fcHMRF as this fully connected CRF to appropriately model the complex spatial dependencies in neuroimaging data.
Unlike standard fully connected CRF applications in supervised image segmentation, where large numbers of labeled images are available for training, our multiple testing context is unsupervised. We do not observe the ground-truth hypothesis states and are limited to a single image of test statistics and their transformations. To address this, we develop an expectation-maximization (EM) algorithm to estimate the parameters of the fcHMRF model and subsequently compute the LIS values for hypothesis inference. Since exact inference in the fcHMRF is computationally intractable, we employ the mean-field approximation method [35, 32] within both the EM algorithm and LIS computation. To implement the mean-field updates efficiently, we leverage the CRFs-as-Recurrent-Neural-Networks (CRF-RNN) technique [33], which unrolls mean-field iterations into a series of differentiable operations analogous to layers in a neural network. In addition, we incorporate the permutohedral lattice filtering technique [36, 37] to accelerate the message-passing step within each mean-field iteration. Together, these techniques reduce the time complexity from quadratic to linear in the number of tests, enabling scalable inference for large neuroimaging datasets. The resulting fcHMRF-LIS method thus combines spatial expressiveness, statistical stability, and computational efficiency.
Our main contributions are summarized as follows:
  • We propose fcHMRF-LIS, a novel spatial FDR control method for voxel-wise multiple testing in neuroimaging data. The method is built upon our newly designed fcHMRF, which effectively captures complex spatial dependencies, including spatial heterogeneity, distance-related dependence, and long-range interactions, all of which are not jointly addressed by existing methods.
  • The proposed fcHMRF-LIS integrates the LIS-based testing procedure [26] to control the FDR while minimizing the FNR. The underlying fcHMRF model is specifically designed with a small set of parameters, enabling fcHMRF-LIS to maintain low variability in both FDP and FNP across replications, thereby offering improved stability over current deep learning and local smoothing-based methods.
  • We develop an efficient EM algorithm for estimating the parameters of the fcHMRF model. To enable scalable inference, we incorporate the mean-field approximation [35, 32], CRF-RNN technique [33], and permutohedral lattice filtering [36, 37], reducing the time complexity from quadratic to linear in the number of tests.
  • We conduct extensive simulation studies to compare fcHMRF-LIS with eight existing FDR control methods [9, 10, 11, 12, 20, 17, 18, 21]. The results demonstrate the superiority of fcHMRF-LIS in accurate FDR control, lower FNR, higher TP, and reduced variability across replications.
  • We apply fcHMRF-LIS to a large-scale FDG-PET dataset from the Alzheimer’s Disease Neuroimaging Initiative (ADNI), analyzing 439,758 brain voxels to identify regions associated with metabolic changes across AD-related stages. The fcHMRF-LIS yields neurobiologically relevant findings and achieves significantly improved computational efficiency, requiring only about 1.2 hours on a 20-core CPU server, with no GPU. In contrast, the nnHMRF-LIS method [12] requires about 5.6 hours, even with GPU acceleration.
The rest of the paper is organized as follows. Section 2 introduces the proposed fcHMRF-LIS method and its mean-field EM algorithm for parameter estimation. Section 3 compares fcHMRF-LIS with eight existing FDR control methods via extensive simulations. Section 4 applies these nine methods to analyze the FDG-PET data from ADNI. Section 5 concludes the paper. Detailed implementation of the nine methods is provided in A. A Python package implementing the fcHMRF-LIS method is publicly available at https://github.com/kimtae55/fcHMRF-LIS.
2 Proposed Method
2.1 Problem Formulation
Consider the problem of testing voxel-wise differences in population means between two comparison groups based on a given measurement. In our motivating AD study, to compare the brain glucose metabolism between two groups, for example, the AD group and the cognitively normal group, we test voxel-wise differences in their population means of the standardized uptake value ratio (SUVR) obtained from brain FDG-PET images. Assume each subject has a 3D brain image with $m$ voxels of interest. Let $x_{i}$ be the test statistic for the null hypothesis $\mathcal{H}_{i}$, which states no difference in the mean SUVR values between the two groups at voxel $i$, i.e., $\mu_{i1}-\mu_{i2}=0$. Define $h_{i}$ as a hidden state representing the unobservable truth of $\mathcal{H}_{i}$, where $h_{i}=1$ if $\mathcal{H}_{i}$ is false, and $h_{i}=0$ if $\mathcal{H}_{i}$ is true. The goal of multiple testing is to infer the unknown hypothesis states $\boldsymbol{h}=(h_{1},\dots,h_{m})$ from the observed test statistics $\boldsymbol{x}=(x_{1},\dots,x_{m})$.
We aim to develop an optimal FDR control method for voxel-wise multiple testing, while also maintaining low variability in both FDP and FNP. The FDR and FNR are the means of FDP and FNP, respectively, formulated as
$\text{FDR}=\mathbb{E}\big{[}\text{FDP}\big{]}\quad$ $\text{and}\quad\text{FNR}=\mathbb{E}\big{[}\text{FNP}\big{]},\quad\text{with}$
$\text{FDP}=\frac{N_{10}}{\max(N_{1},1)}\quad$ $\text{and}\quad\text{FNP}=\frac{N_{01}}{\max(N_{0},1)},$
where the classification notation used for tested hypotheses is summarized in Table 1. Given a prespecified FDR level $\alpha$, an FDR control method is valid if it controls the FDR at level $\alpha$, and is optimal if it minimizes the FNR among all valid FDR control methods. For simplicity, false nulls and rejected nulls are referred to as signals and discoveries, respectively.
Table 1: Classification of tested hypotheses
Number Not rejected Rejected Total
True null $N_{00}$ $N_{10}$ $m_{0}$
False null $N_{01}$ $N_{11}$ $m_{1}$
Total $N_{0}$ $N_{1}$ $m$
2.2 LIS-based Testing Procedure
The local index of significance (LIS) was proposed in [26] to replace the p-value for multiple testing under dependence. The LIS for hypothesis $\mathcal{H}_{i}$ is defined by
$\operatorname{LIS}_{i}(\boldsymbol{x})=p(h_{i}=0|\boldsymbol{x}),$ (1)
which is the conditional probability of $h_{i}=0$ given all test statistics $\boldsymbol{x}=(x_{1},...,x_{m})$ not just the local statistic $x_{i}$. The LIS-based testing procedure for controlling the FDR at a prespecified level $\alpha$ is given as follows:
$\begin{split}&\mbox{Let }k=\max\{j:\frac{1}{j}\sum_{i=1}^{j}\operatorname{LIS}_{(i)}(\boldsymbol{x})\leq\alpha\},\\ &\mbox{then reject all }\mathcal{H}_{(i)}\mbox{ with }i=1,...,k.\end{split}$ (2)
Here, $\operatorname{LIS}_{(i)}(\boldsymbol{x})$ is the $i$-th smallest LIS value and $\mathcal{H}_{(i)}$ is the corresponding null hypothesis. Due to the identity
$\text{FDR}=\mathbb{E}\left[\frac{1}{\max(R,1)}\sum_{i=1}^{R}\operatorname{LIS}_{(i)}(\boldsymbol{x})\right],$
the LIS-based testing procedure in (LABEL:def:_LIS_procedure) is valid for controlling FDR at level $\alpha$. Under mild conditions, it is also asymptotically optimal for minimizing the FNR [26, 27, 12].
In practice, the true values $\{\operatorname{LIS}_{i}(\boldsymbol{x})\}_{i=1}^{m}$ are unknown and must be replaced by estimates $\{\widehat{\operatorname{LIS}}_{i}(\boldsymbol{x})\}_{i=1}^{m}$ obtained from a fitted model for the joint distribution $p(\boldsymbol{x},\boldsymbol{h})$. The effectiveness of the LIS-based testing procedure thus hinges on how accurately the selected model captures the dependence among tests. To model the complex spatial dependencies present in neuroimaging data, we propose an fcHMRF as the working model for $p(\boldsymbol{x},\boldsymbol{h})$. This fcHMRF generalizes the nearest-neighbor HMRF considered by [12]. Since the validity and optimality proofs in [12] apply to a broad class of HMRF models, including our proposed fcHMRF, the LIS-based testing procedure using the fcHMRF retains the validity and optimality.
2.3 The fcHMRF Model
We propose a novel fcHMRF to effectively capture the spatial dependence among tests. The proposed fcHMRF assumes the observed test statistics $\boldsymbol{x}=(x_{1},\dots,x_{m})$, are conditionally independent given the hidden states $\boldsymbol{h}=(h_{1},\dots,h_{m})$, leading to the conditional probability density function:
$\begin{split}p(\boldsymbol{x}|\boldsymbol{h})&=\prod_{i=1}^{m}p(x_{i}|h_{i})\quad\text{with}\\ p(x_{i}|h_{i})&=(1-h_{i})f_{0}(x_{i})+h_{i}f_{1}(x_{i}),\end{split}$ (3)
where $f_{0}$ is a known null density function and $f_{1}$ is an unknown non-null density function. In our context of testing the equality of two means, we define each $x_{i}$ as a z-statistic which can be easily converted from a t-statistic [12]. Thus, we set the null density $f_{0}$ to be the standard Gaussian density:
$f_{0}(x)=\varphi(x)=\frac{1}{\sqrt{2\pi}}\exp(-\frac{x^{2}}{2}).$ (4)
The unknown non-null density $f_{1}$ will be estimated nonparametrically using kernel density estimation.
Our fcHMRF models the probability of hidden states $\boldsymbol{h}$ by a fully connected Markov random field as
$p(\boldsymbol{h})=\frac{1}{Z}\exp\Big{(}-w_{0}\sum_{i=1}^{m}h_{i}-\sum_{1\leq i<j\leq m}w_{ij}|h_{i}-h_{j}|\Big{)},$ (5)
where $Z$ is the normalizing constant, and $w_{0}$ and $w_{ij}\neq 0$ are parameters for unary and pairwise potentials, respectively. In particular, $w_{ij}$ modulates the penalty for assigning different hidden state values to voxels $i,j$ based on their properties. Since $w_{ij}\neq 0$ for all pairs of voxels, the hidden states $h_{i}$ and $h_{j}$ of any two voxels are conditionally dependent given the hidden states of all other voxels. Thus, every pair $(h_{i},h_{j})$ is connected in the undirected graph representing the Markov random field $p(\boldsymbol{h})$ [35], making it a fully connected graph that enables long-range interactions.
We model $w_{ij}$ as a combination of the appearance and smoothness kernels:
$w_{ij}$ $=w_{1}\underbrace{\exp(-\sum_{s\in S}\frac{|\ell_{i,s}-\ell_{j,s}|^{2}}{2\theta_{\alpha,s}^{2}}-\frac{|\Delta\mu_{i}-\Delta\mu_{j}|^{2}}{2\theta_{\beta}^{2}})}_{\text{\normalsize appearance kernel}}$ (6)
$\qquad~{}~{}~{}~{}~{}+w_{2}\underbrace{\exp(-\sum_{s\in S}\frac{|\ell_{i,s}-\ell_{j,s}|^{2}}{2\theta_{\gamma,s}^{2}})}_{\text{\normalsize smoothnesskernel}}.$
where $\ell_{i,s}$, $s\in S=\{x,y,z\}$, are the 3D coordinates of voxel $i$, and $\Delta\mu_{i}$ is the difference in mean values at voxel $i$ between the two groups, which is estimated by $\Delta\mu_{i}\approx\Delta\hat{\mu}_{i}=\bar{v}_{i1}-\bar{v}_{i2},$ with $\bar{v}_{ig}$ being the sample mean at voxel $i$ in group $g\in\{1,2\}$. Note that in the appearance kernel, the difference $\Delta\mu_{i}-\Delta\mu_{j}$ matters more than the individual values $\Delta\mu_{i}$ and $\Delta\mu_{j}$.
The above kernel combination is adapted from the one used in fully connected CRFs in image segmentation [32]. In our spatial multiple testing context, we replace the color intensity at voxel $i$ used in their segmentation framework with the mean difference $\Delta\mu_{i}$. In image segmentation, the appearance kernel is inspired by the observation that nearby voxels with similar colors ($\Delta\mu_{i}$ in our context) are likely to have the same class (hypothesis state in our context), and the smoothness kernel removes small isolated regions. The degrees of nearness and similarity are controlled by $\{\theta_{\alpha,s},\theta_{\gamma,s}\}_{s\in S}$ and $\theta_{\beta}$, respectively. The two kernels, together with the fully connected graphical model, can effectively capture spatial heterogeneity, distance-related dependence, and long-range interactions [32, 33, 34].
To simplify the model parameter estimation, we fix $\{\theta_{\alpha,s},\theta_{\gamma,s}\}$ and $\theta_{\beta}$ to be the SDs of $\{\ell_{i,s}-\ell_{j,s}\}_{i\neq j}$ and $\{\Delta\mu_{i}-\Delta\mu_{j}\}_{i\neq j}$, respectively, because these $\theta$ parameters can be approximately absorbed into the corresponding $w$ parameters. Specifically, let $\theta\in\{\theta_{\alpha,s},\theta_{\gamma,s},\theta_{\beta}\}$, $\delta_{ij}\in\{\ell_{i,s}-\ell_{j,s},\Delta\mu_{i}-\Delta\mu_{j}\}$, and $w\in\{w_{1},w_{2}\}$. We use the ordinary least squares (OLS) to approximate $\exp\{-\delta_{ij}^{2}/(2\theta^{2})\}\in(0,1]$ as
$\exp\left(-\frac{\delta_{ij}^{2}}{2\theta^{2}}\right)\approx b(\theta,\{\delta_{ij}\}_{i\neq j})\cdot\exp\left(-\frac{\delta_{ij}^{2}}{2\operatorname{SD}^{2}(\{\delta_{ij}\}_{i\neq j})}\right).$
Then, the OLS coefficient $b(\theta,\{\delta_{ij}\}_{i\neq j})$ can be absorbed into the corresponding learnable parameter $w$. For very large $m$, to avoid the computational burden of exactly computing $\operatorname{SD}(\{\delta_{ij}\}_{i\neq j})$, which requires calculating $\{\delta_{ij}\}_{i\neq j}$ with $O(m^{2})$ time complexity, we instead estimate it using a large yet computationally manageable subsample of $\{\delta_{ij}\}_{i\neq j}$.
From equations (3)–(5), we obtain the conditional probability of $\boldsymbol{h}$ given $\boldsymbol{x}$:
$p(\boldsymbol{h}|\boldsymbol{x})\propto\exp\Bigg{\{}-\sum_{i=1}^{m}h_{i}\Big{[}w_{0}+\log\frac{\varphi(x_{i})}{f_{1}(x_{i})}\Big{]}-\sum_{1\leq i<j\leq m}w_{ij}|h_{i}-h_{j}|\Bigg{\}},$ (7)
which also represents a fully connected Markov random field.
The parameter set of the fcHMRF is $\boldsymbol{\phi}=\{f_{1},\boldsymbol{w}=(w_{0},w_{1},w_{2})\}$. From the definition in (1), $\operatorname{LIS}_{i}(\boldsymbol{x})=p(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi})$ is the marginal probability that $h_{i}=0$ under the conditional distribution $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})$. Computing the exact marginal probability $p(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi})$ is intractable, as it requires summing over all $2^{m}$ possible configurations of $\boldsymbol{h}$ to obtain the normalizing constant in $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})$, followed by $2^{m-1}$ addition operations to marginalize out $\boldsymbol{h}_{-i}$ (all entries of $\boldsymbol{h}$ except $h_{i}$). One might consider approximating $\operatorname{LIS}_{i}(\boldsymbol{x})$ using the Gibbs sampler by $\operatorname{LIS}_{i}(\boldsymbol{x})\approx N^{-1}\sum_{n=1}^{N}\mathbb{I}(h_{i}^{(n,\boldsymbol{x})}=0)$ [12], where $\{\boldsymbol{h}^{(n,\boldsymbol{x})}=(h_{1}^{(n,\boldsymbol{x})},\dots,h_{m}^{(n,\boldsymbol{x})})\}_{n=1}^{N}$ are $N$ samples generated by the Gibbs sampler [38] from $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})$, and $\mathbb{I}(\cdot)$ is the indicator function. However, this approximation approach is computationally expensive due to the time complexity $O(m^{2}(B+N))$ for the fully connected Markov random field $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})$, where $B$ is the number of burn-in iterations. Instead, we approximate $\operatorname{LIS}_{i}(\boldsymbol{x})$ by
$\operatorname{LIS}_{i}(\boldsymbol{x})=p(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi})\approx q_{i}(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi}),$ (8)
using the mean-field approximation $q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})=\prod_{i=1}^{m}q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi})$ to $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi})$ [32]. The mean-field approximation is efficiently computed using the permutohedral lattice filtering method [36, 37], which reduces the time complexity to $O(mR)$, where $R$ is the number of mean-field iterations. The details on the mean-field approximation and the permutohedral lattice filtering method are provided in Sections 2.5 and 2.6, respectively.
2.4 The EM Algorithm for the Estimation of fcHMRF
In practice, the parameter set $\boldsymbol{\phi}$ of the fcHMRF is unknown and is replaced with its estimate when computing the LIS values using equation (8). We develop an EM algorithm to estimate $\boldsymbol{\phi}$.
Let $\boldsymbol{\phi}^{(t+1)}$ denote the final estimate of $\boldsymbol{\phi}$ in the $(t+1)$-th iteration of the EM algorithm, which aims to maximize the following objective function
$Q(\boldsymbol{\phi}|\boldsymbol{\phi}^{(t)})=\sum_{\boldsymbol{h}}p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log p(\boldsymbol{x},\boldsymbol{h};\boldsymbol{\phi})$ (9)
$=\underbrace{\sum_{\boldsymbol{h}}p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log p\bigl{(}\boldsymbol{x}|\boldsymbol{h};f_{1}\bigr{)}}_{\mbox{\normalsize$Q_{1}(f_{1}|\boldsymbol{\phi}^{(t)})$}}+\underbrace{\sum_{\boldsymbol{h}}p\bigl{(}\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)}\bigr{)}\log p\bigl{(}\boldsymbol{h};\boldsymbol{w}\bigr{)}}_{\mbox{\normalsize$Q_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$}}.$
We can maximize $Q(\boldsymbol{\phi}|\boldsymbol{\phi}^{(t)})$ for $\boldsymbol{\phi}$ by maximizing $Q_{1}(f_{1}|\boldsymbol{\phi}^{(t)})$ for $f_{1}$ and $Q_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ for $\boldsymbol{w}$, separately.
However, similar to computing $\{\operatorname{LIS}_{i}(\boldsymbol{x})\}_{i=1}^{m}$ as mentioned at the end of Section 2.3, the exact maximization of $Q(\boldsymbol{\phi}|\boldsymbol{\phi}^{(t)})$ is intractable due to the normalizing constants in fully connected Markov random fields $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $p(\boldsymbol{h};\boldsymbol{w})$, and the approximate maximization using Gibbs-sampler samples from these fields is computationally expensive because each update of $\boldsymbol{\phi}$ within the $(t+1)$-th EM iteration requires $O(m^{2}(B+N))$ time [12]. To address this, following [39], we propose using the mean-field approximations of $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $p(\boldsymbol{h};\boldsymbol{w})$ in the EM algorithm, but it still costs $O(m^{2}R+mN)$ time per update based on $N$ mean-field samples. We thus accelerate it using the permutohedral lattice filtering method [36, 37], reducing the time complexity to $O(m(R+N))$ per update (see Section 2.6).
In $Q_{1}(f_{1}|\boldsymbol{\phi}^{(t)})$, replacing $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ with its mean-field approximation $q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})=\prod_{i=1}^{m}q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ yields
$Q_{1}(f_{1}|\boldsymbol{\phi}^{(t)})\approx\hat{Q}_{1}(f_{1}|\boldsymbol{\phi}^{(t)})$
$=\sum_{\boldsymbol{h}}q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log p(\boldsymbol{x}|\boldsymbol{h};f_{1})$
$=\sum_{\boldsymbol{h}}q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\sum_{i}[h_{i}\log f_{1}(x_{i})+(1-h_{i})\log\varphi(x_{i})]$
$=\sum_{i}\sum_{\boldsymbol{h}}[h_{i}q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})]\log f_{1}(x_{i})$
$\qquad+\sum_{i}\sum_{\boldsymbol{h}}[(1-h_{i})q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})]\log\varphi(x_{i})$
$=\sum_{i}q_{i}(h_{i}=1|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log f_{1}(x_{i})$
$\qquad+\sum_{i}q_{i}(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log\varphi(x_{i}).$ (10)
The optimal Gaussian-kernel estimator [28] of $f_{1}$ for maximizing $\hat{Q}_{1}(f_{1}|\boldsymbol{\phi}^{(t)})$ is
$f_{1}^{(t+1)}(x)=\frac{\sum_{i=1}^{m}q_{i}(h_{i}=1|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\frac{1}{h}\varphi(\frac{x-x_{i}}{h})}{\sum_{i=1}^{m}q_{i}(h_{i}=1|\boldsymbol{x};\boldsymbol{\phi}^{(t)})}.$ (11)
We use the rule-of-thumb bandwidth [40]:
$h=0.9\cdot\min\left\{\operatorname{SD}(\{x_{i}\}_{i=1}^{m}),\frac{\text{IQR}(\{x_{i}\}_{i=1}^{m})}{1.34}\right\}\cdot m_{\text{eff}}^{-1/5},$
where $\text{IQR}(\{x_{i}\}_{i=1}^{m})$ is the interquartile range of $\{x_{i}\}_{i=1}^{m}$, and the effective sample size [41]
$m_{\text{eff}}=\frac{[\sum_{i=1}^{m}q_{i}(h_{i}=1|\boldsymbol{x};\boldsymbol{\phi}^{(t)})]^{2}}{\sum_{i=1}^{m}[q_{i}(h_{i}=1|\boldsymbol{x};\boldsymbol{\phi}^{(t)})]^{2}}.$
In $Q_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$, replacing $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $p(\boldsymbol{h};\boldsymbol{w})$ with their mean-field approximations $q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})=\prod_{i=1}^{m}q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $q(\boldsymbol{h};\boldsymbol{w})=\prod_{i=1}^{m}q_{i}(h_{i};\boldsymbol{w})$, and then applying the Monte Carlo approximation, we obtain
$Q_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ $\approx\sum_{\boldsymbol{h}}q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\log q(\boldsymbol{h};\boldsymbol{w})$
$\approx\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})=\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{m}\log q_{i}(h_{i}=h_{i}^{(n,t)};\boldsymbol{w}),$ (12)
where $h_{i}^{(n,t)}$ is independently sampled from the Bernoulli distribution $q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ for all $n=1,\dots,N$ and $i=1,\dots,m$.
The optimal $\boldsymbol{w}$ that minimizes $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ is obtained using gradient descent, where the gradient is efficiently computed via backpropagation [42], a reverse-mode automatic differentiation method, through the CRF-RNN network [33] that implements the mean-field approximation $\{q_{i}(h_{i};\boldsymbol{w})\}_{i=1}^{m}$. The mean-field approximation and its CRF-RNN implementation are described in Section 2.5.
2.5 Mean-field Approximation and CRF-RNN
In the EM algorithm, we replace the Markov random fields $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $p(\boldsymbol{h};\boldsymbol{w})$ in (7) and (5) with their mean-field approximations $q(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})=\prod_{i=1}^{m}q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $q(\boldsymbol{h};\boldsymbol{w})=\prod_{i=1}^{m}q_{i}(h_{i};\boldsymbol{w})$. We write $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and $p(\boldsymbol{h};\boldsymbol{w})$ in the form of CRFs, which are widely used in image segmentation [32, 33, 34]:
$p(\boldsymbol{h}|\boldsymbol{I})\,{=}\,\frac{1}{Z(\boldsymbol{I})}\exp\left\{\sum_{i=1}^{m}U_{i}(h_{i})\,{-}\,\sum_{1\leq i<j\leq m}|h_{i}\,{-}\,h_{j}|\sum_{l=1}^{2}w_{l}k_{l}(\boldsymbol{I}_{i},\boldsymbol{I}_{j})\right\}$ (13)
where image features $\boldsymbol{I}=\{\boldsymbol{I}_{i}\}_{i=1}^{m}$ with $\boldsymbol{I}_{i}=\{\{\ell_{i,s}\}_{s\in S},\Delta\mu_{i}\}$, and $k_{l}(\cdot,\cdot)$ is a Gaussian kernel. For simplicity, we omit the conditioning on $\boldsymbol{x}$, if applicable, in the notation $p(\boldsymbol{h}|\boldsymbol{I})$. We also drop the conditioning on $\boldsymbol{I}$ in $p(\boldsymbol{h}|\boldsymbol{I})$ throughout the rest of the paper.
The mean-field method [35] finds a distribution $q(\boldsymbol{h})$, which is close to $p(\boldsymbol{h})$ in the sense that it minimizes the Kullback–Leibler divergence
$D_{\text{KL}}(q(\boldsymbol{h})\|p(\boldsymbol{h}))=\sum_{\boldsymbol{h}}q(\boldsymbol{h})\log\frac{q(\boldsymbol{h})}{p(\boldsymbol{h})}$
within the class of distributions representable as a product of independent marginals:
$q(\boldsymbol{h})=\prod_{i=1}^{m}q_{i}(h_{i}).$
The mean-field approximation $q(\boldsymbol{h})$ for $p(\boldsymbol{h})$ given in (13) can be iteratively obtained by the following update equation [32]:
$q_{i}(h)=\frac{1}{Z_{i}}\exp\left\{U_{i}(h)-\sum_{h^{\prime}}|h-h^{\prime}|\sum_{l=1}^{2}w_{l}\sum_{j\neq i}k_{l}(\boldsymbol{I}_{i},\boldsymbol{I}_{j})q_{j}(h^{\prime})\right\}.$
The CRF-RNN method [33] decomposes a single mean-field iteration from the above update equation into simpler steps that can be implemented using neural network operations, as shown in Algorithm 1. CRF-RNN further formulates the mean-field approximation algorithm, consisting of multiple mean-field iterations, as an RNN [42], a type of neural network, where the output of one iteration is used as the input of the next iteration and all iterations share the same parameters. This makes implementing mean-field approximation simple and efficient by leveraging standard neural network libraries such as PyTorch [43] and TensorFlow [44]. The CRF-RNN formulation of mean-field iterations is given by the following equations:
$\boldsymbol{H}_{1}(r)$ =$=\begin{cases}\text{softmax}(\boldsymbol{U}),&r=0,\\ \boldsymbol{H}_{2}(r-1),&0<r\leq R,\end{cases}$ (14)
$\boldsymbol{H}_{2}(r)$ $=\boldsymbol{f}_{\boldsymbol{\psi}}(\boldsymbol{U},\boldsymbol{H}_{1}(r),\boldsymbol{I}),\quad 0\leq r\leq R,$ (15)
$\boldsymbol{Y}(r)$ =$=\begin{cases}\boldsymbol{0},&0\leq r<R,\\ \boldsymbol{H}_{2}(r),&r=R,\end{cases}$ (16)
where $\boldsymbol{U}=\{U_{i}(h)\}_{i=1}^{m}$ is the set of voxel-wise unary potential values, $\boldsymbol{f}_{\boldsymbol{\psi}}(\boldsymbol{U},\boldsymbol{q}_{\text{in}},\boldsymbol{I})$ is the estimation of marginal distributions $\{q_{i}\}_{i=1}^{m}$ after one mean-field iteration, $\boldsymbol{q}_{\text{in}}$ is the estimation from the previous iteration, and $R$ is the number of iterations. The parameter set $\boldsymbol{\psi}$ is $\boldsymbol{\phi}^{(t)}$ for $p(\boldsymbol{h}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ and is $\boldsymbol{w}$ for $p(\boldsymbol{h};\boldsymbol{w})$. Figure 2 shows the CRF-RNN network. The key advantage of CRF-RNN is that the parameter set $\boldsymbol{\psi}$ can be learned using backpropagation [42], a reverse-mode automatic differentiation method for neural networks, available in PyTorch and TensorFlow, to efficiently compute the gradient for gradient descent.
[…]
1:$q_{i}(h)\leftarrow\dfrac{\exp(U_{i}(h))}{\sum_{h^{\prime}\in\{0,1\}}\exp(U_{i}(h^{\prime}))},~{}i=1,\dots,m$triangleright Initialization
2:repeat
3:     $\tilde{q}_{i,l}(h)\leftarrow\sum_{j\neq i}k_{l}(\boldsymbol{I}_{i},\boldsymbol{I}_{j})q_{j}(h),~{}i=1,\dots,m$
4$\triangleright$ Message passing
5:     for $i=1,\dots,m$ do
6:         $\check{q}_{i}(h)\leftarrow\sum_{l=1}^{2}w_{l}\tilde{q}_{i,l}(h)$triangleright Weighting filter outputs
7:         $\hat{q}_{i}(h)\leftarrow\sum_{h^{\prime}\in\{0,1\}}|h-h^{\prime}|\check{q}_{i}(h^{\prime})$
8$\triangleright$ Compatibility transform
9:         $\breve{q}_{i}(h)\leftarrow U_{i}(h)-\hat{q}_{i}(h)$triangleright Add unary potentials
10:         $q_{i}(h)\leftarrow\dfrac{\exp(\breve{q}_{i}(h))}{\sum_{h^{\prime}\in\{0,1\}}\exp(\breve{q}_{i}(h^{\prime}))}$triangleright Normalizing
11:     end for
12:until stopping criterion met
Figure 2: The CRF-RNN network. Gating functions $G_{1}$ and $G_{2}$ are given in (14) and (16), respectively.
2.6 Permutohedral Lattice Filtering
The message-passing step in Algorithm 1 for mean-field approximation can be formulated as Gaussian filtering:
$v_{i}^{\prime}=\sum_{j=1}^{m}\exp\left(-\frac{1}{2}\|\boldsymbol{p}_{i}-\boldsymbol{p}_{j}\|_{2}^{2}\right)v_{j},~{}~{}i=1,\dots,m,$ (17)
where the input value $v_{j}=q_{j}(h)$, the output value $v_{i}^{\prime}=\tilde{q}_{i,l}(h)+q_{i}(h)$, and the position vector $\boldsymbol{p}_{i}\in\mathbb{R}^{d}$ is $\big{(}(\ell_{i,s}/\theta_{\alpha,s})_{s\in S},\Delta\mu_{i}/\theta_{\beta}\big{)}^{\top}$ for $l=1$ and is $(\ell_{i,s}/\theta_{\gamma,s})_{s\in S}^{\top}$ for $l=2$ in our context.
The exact computation of this Gaussian filtering is computationally expensive, with time complexity $O(m^{2}d)$. To address this, the permutohedral lattice filtering method [36, 37] offers a highly efficient approximation with time complexity $O(md^{2})$ and space complexity $O(md)$. This method decomposes the filtering operation into three stages named splatting, blurring and slicing, which correspond to the familiar notions of embedding, convolution, and reconstruction in machine learning.
Let $H_{d}=\{\boldsymbol{s}\in\mathbb{R}^{d+1}|\boldsymbol{s}^{\top}\boldsymbol{1}_{d+1}=0\}$ be the $d$-dimensional hyperplane consisting of vectors in $\mathbb{R}^{d+1}$ whose coordinates sum to zero, where $\boldsymbol{1}_{d+1}$ is the $(d+1)$-dimensional vector of all ones. The permutohedral lattice in $H_{d}$, denoted by $A_{d}^{*}$, is defined as the projection of the scaled regular grid $(d+1)\mathbb{Z}^{d+1}$ onto $H_{d}$. The scale factor $d+1$ ensures that lattice coordinates remain integer-valued. Equivalently, the permutohedral lattice $A_{d}^{*}$ can be expressed as the union of all remainder-$k$ points in $H_{d}$ for $k=0,\dots,d$:
$A_{d}^{*}=\bigcup_{k=0}^{d}\left\{\boldsymbol{s}\in H_{d}\,|\,\boldsymbol{s}=(d+1)\boldsymbol{z}+k\boldsymbol{1}_{d+1},\boldsymbol{z}\in\mathbb{Z}^{d+1}\right\}.$
The permutohedral lattice $A_{d}^{*}$ tessellates the hyperplane $H_{d}$ with uniform simplices, each of which is a permutation and translation of a canonical simplex whose vertices are given by
$\boldsymbol{s}_{k}=(k\boldsymbol{1}_{d+1-k}^{\top},-(d+1-k)\boldsymbol{1}_{k}^{\top})^{\top},~{}~{}k=0,\dots,d,$
where each $\boldsymbol{s}_{k}$ is a remainder-$k$ point.
In the splatting stage, the input values $\{v_{i}\}_{i=1}^{m}$ are embedded onto the permutohedral lattice $A_{d}^{*}$. Specifically, the position vector $\boldsymbol{p}_{i}$ of value $v_{i}$ is embedded into $H_{d}$ as $\mathbf{E}\boldsymbol{p}_{i}$, where $\mathbf{E}\in\mathbb{R}^{(d+1)\times d}$ has orthonormal columns spanning $H_{d}$:
$\mathbf{E}=\left(\begin{array}[]{cccc}1&1&\cdots&1\\ -1&1&\cdots&1\\ 0&-2&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-d\end{array}\right)\left(\begin{array}[]{cccc}\frac{1}{\sqrt{2}}&0&\cdots&0\\ 0&\frac{1}{\sqrt{6}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\frac{1}{\sqrt{d(d+1)}}\end{array}\right).$
The embedded point $\mathbf{E}\boldsymbol{p}_{i}$ lies inside a simplex in $A_{d}^{*}$, whose vertices are denoted by $\{\boldsymbol{s}_{i,k}\}_{i=0}^{d}$. Let $\boldsymbol{b}_{i}=(b_{i,0},\dots,b_{i,d})$ denote the barycentric coordinates of $\mathbf{E}\boldsymbol{p}_{i}$ within this simplex, so that
[…]
Then the value $v_{i}$ of point $\mathbf{E}\boldsymbol{p}_{i}$ is distributed to each vertex $\boldsymbol{s}_{i,k}$ with weight $b_{i,k}$, i.e., $b_{i,k}v_{i}$ is accumulated at $\boldsymbol{s}_{i,k}$.
In the blurring stage, the embedded values on the permutohedral lattice $A_{d}^{*}$ are smoothed by a convolution that approximates Gaussian filtering. This is implemented by convolving with the kernel $(1,2,1)$ along each of the $d+1$ lattice directions of the form ±$\pm(1,\dots,1,-d,1,\dots,1)$.
The slicing stage mirrors the splatting stage in reverse. Each point $\{\mathbf{E}\boldsymbol{p}_{i}\}_{i=1}^{m}$ gathers the blurred values from the vertices of its enclosing simplex in lattice $A_{d}^{*}$ using the same barycentric weights $\boldsymbol{b}_{i}$, thereby reconstructing the filtered output $\{v_{i}^{\prime}\}_{i=1}^{m}$.
The permutohedral lattice filtering method is summarized in Algorithm 2.
[…]
1:Input values $\{v_{i}\}_{i=1}^{m}$ and position vectors $\{\boldsymbol{p}_{i}\}_{i=1}^{m}$
2:$V(\boldsymbol{s})\leftarrow 0,~{}\forall\boldsymbol{s}\in A_{d}^{*}$
3:for $i=1,\dots,m$ do $\triangleright$ Splatting
4:     for $k=0,\dots,d$ do
5:         $\boldsymbol{s}\leftarrow$ the closest remainder-$k$ point in $A_{d}^{*}$ to $\mathbf{E}\boldsymbol{p}_{i}$ (by Theorem 4.10 in [37])
6:         $b\leftarrow$ the barycentric coordinate of $\mathbf{E}\boldsymbol{p}_{i}$ at $\boldsymbol{s}$ (by Proposition 5.2 in [37])
7:         $V(\boldsymbol{s})\leftarrow V(\boldsymbol{s})+bv_{i}$
8:     end for
9:end for
10:for $k=0,\dots,d$ do $\triangleright$ Blurring
11:     $\boldsymbol{t}\leftarrow(\boldsymbol{1}_{k}^{\top},-d,\boldsymbol{1}_{d-k}^{\top})^{\top}$
12:     for all $\boldsymbol{s}\in A_{d}^{*}$ do
13:         $W(\boldsymbol{s})\leftarrow\dfrac{V(\boldsymbol{s}-\boldsymbol{t})}{4}+\dfrac{V(\boldsymbol{s})}{2}+\dfrac{V(\boldsymbol{s}+\boldsymbol{t})}{4}$
14:     end for
15:     $V\leftarrow W$
16:end for
17:for $i=1,\dots,m$ do $\triangleright$ Slicing
18:     $v_{i}^{\prime}\leftarrow 0$
19:     for $k=0,\dots,d$ do
20:         $\boldsymbol{s}\leftarrow$ the closest remainder-$k$ point in $A_{d}^{*}$ to $\mathbf{E}\boldsymbol{p}_{i}$
21:         $b\leftarrow$ the barycentric coordinate of $\mathbf{E}\boldsymbol{p}_{i}$ at $\boldsymbol{s}$
22:         $v_{i}^{\prime}\leftarrow v_{i}^{\prime}+bW(\boldsymbol{s})$
23:     end for
24:end for
25:$\{v_{i}^{\prime}\}_{i=1}^{m}$
2.7 The Complete fcHMRF-LIS Algorithm
From equation (8), the LIS estimates are given by $\widehat{\text{LIS}}_{i}(\boldsymbol{x})=q_{i}(h_{i}=0|\boldsymbol{x};\hat{\boldsymbol{\phi}})$ for $i=1,\dots,m$, where $\prod_{i=1}^{m}q_{i}(h_{i}|\boldsymbol{x};\hat{\boldsymbol{\phi}})$ is the mean-field approximation of $p(\boldsymbol{h}|\boldsymbol{x};\hat{\boldsymbol{\phi}})$ under the proposed fcHMRF model in Section 2.3, with $\hat{\boldsymbol{\phi}}$ denoting the estimate of the model parameter set $\boldsymbol{\phi}=\{f_{1},\boldsymbol{w}\}$. To obtain $\hat{\boldsymbol{\phi}}$, we apply the mean-field EM algorithm described in Section 2.4. In the $(t+1)$-th EM iteration, the non-null density estimate $f_{1}^{(t+1)}$ is computed using equation (11), and the weight estimate $\boldsymbol{w}^{(t+1)}$ is obtained by minimizing $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ in (2.4) via gradient descent. We implement gradient descent using the AdamW optimizer [45], with the gradient of $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ is automatically computed via backpropagation [42]. The mean field approximations $\{q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\}_{i=1}^{m}$ and $\{q_{i}(h_{i};\boldsymbol{w})\}_{i=1}^{m}$, used in (11) and (2.4), are computed using the mean-field algorithm in Algorithm 1, efficiently implemented via CRF-RNN (14)-(16). The message passing step in Algorithm 1 is approximated using the permutohedral lattice filtering method described in Algorithm 2. We independently sample $h_{i}^{(n,t)}$ from the Bernoulli distribution $q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ for all $n=1,\dots,N$ and $i=1,\dots,m$, and use these samples to compute $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ and its gradient. convergence of the EM algorithm is monitored using the approximate $Q$-function, $\hat{Q}(\boldsymbol{\phi}^{(t+1)}|\boldsymbol{\phi}^{(t)})=\hat{Q}_{1}(f_{1}^{(t+1)}|\boldsymbol{\phi}^{(t)})+\hat{Q}_{2}(\boldsymbol{w}^{(t+1)}|\boldsymbol{\phi}^{(t)})$, where $\hat{Q}_{1}$ and $\hat{Q}_{2}$ are defined in (10) and (2.4), respectively. After obtaining the final estimate $\hat{\boldsymbol{\phi}}$ from the EM algorithm, the LIS estimates, $\widehat{\text{LIS}}_{i}(\boldsymbol{x})=q_{i}(h_{i}=0|\boldsymbol{x};\hat{\boldsymbol{\phi}}),i=1,\dots,m$, are computed via the same mean-field algorithm based on CRF-RNN and permutohedral lattice filtering. The resulting estimates $\{\widehat{\text{LIS}}_{i}(\boldsymbol{x})\}_{i=1}^{m}$ are then plugged into the LIS-based testing procedure (LABEL:def:_LIS_procedure) to obtain the multiple testing results. Algorithm 3 summarizes the steps of the proposed fcHMRF-LIS algorithm. Details of the implementation used in our simulations and real-data analysis are provided in A.
[…]
1:A 3D volume of z-statistics $\boldsymbol{x}$ and corresponding 3D volume of mean-difference estimates $\{\Delta\hat{\mu}_{i}\}_{i=1}^{m}$.
2:Initialize $\boldsymbol{\phi}^{(0)}$.
3:repeat $t=0,1,\dots$triangleright EM algorithm
4:     Compute $\{q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\}_{i=1}^{m}$ via mean-field iterations (Algorithm 1) implemented by CRF-RNN (14)-(16), where message passing (17) is approximated by permutohedral lattice filtering (Algorithm 2).
5:     Compute $f_{1}^{(t+1)}$ from (11) using $\{q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})\}_{i=1}^{m}$.
6:     Compute $\boldsymbol{w}^{(t+1)}$ by minimizing $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ in (2.4):
7:         Sample $h_{i}^{(n,t)}$ independently from $q_{i}(h_{i}|\boldsymbol{x};\boldsymbol{\phi}^{(t)})$ for all $n=1,\dots,N$ and $i=1,\dots,m$;
8:         $\triangleright$ Gradient descent
9:              Compute $\{q_{i}(h_{i};\boldsymbol{w})\}_{i=1}^{m}$ as in Step 4;
10:              Compute the gradient of $-\hat{Q}_{2}(\boldsymbol{w}|\boldsymbol{\phi}^{(t)})$ via backpropagation using $\{q_{i}(h_{i};\boldsymbol{w})\}_{i=1}^{m}$ and $\{\{h_{i}^{(n,t)}\}_{n=1}^{N}\}_{i=1}^{m}$;
11:              Update $\boldsymbol{w}$ using a gradient descent step;
12:         until stopping criterion met.
13:     $\boldsymbol{\phi}^{(t+1)}\leftarrow\{f_{1}^{(t+1)},\boldsymbol{w}^{(t+1)}\}$.
14:     $\hat{Q}(\boldsymbol{\phi}^{(t+1)}|\boldsymbol{\phi}^{(t)})\leftarrow\hat{Q}_{1}(f_{1}^{(t+1)}|\boldsymbol{\phi}^{(t)})+\hat{Q}_{2}(\boldsymbol{w}^{(t+1)}|\boldsymbol{\phi}^{(t)})$.
15:until stopping criterion met.
16:Compute $\{q_{i}(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi}^{(t+1)})\}_{i=1}^{m}$ as in Step 4.
17:$\widehat{\text{LIS}}_{i}(\boldsymbol{x})\leftarrow q_{i}(h_{i}=0|\boldsymbol{x};\boldsymbol{\phi}^{(t+1)})$, $i=1,\dots,m$.
18:Run the LIS-based testing procedure (LABEL:def:_LIS_procedure) with $\{\widehat{\text{LIS}}_{i}(\boldsymbol{x})\}_{i=1}^{m}$.
19:A 3D volume of estimates for the null hypothesis states $\boldsymbol{h}$.
3 Simulation Studies
3.1 Methods for Comparison
We conduct extensive simulations to compare the proposed fcHMRF-LIS method with eight existing FDR control methods, including BH [9], q-value [10], LocalFDR [11], nnHMRF-LIS [12], NeuralFDR [20], SmoothFDR [17], LAWS [18], and DeepFDR [21]. The BH, q-value, and LocalFDR methods are classic FDR control methods developed for independent tests, while nnHMRF-LIS, NeuralFDR, SmoothFDR, LAWS, and DeepFDR are spatial methods applicable to 3D image data.
(a) The means of FDP, FNP, and TP for the cube with about $10\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.64.
(b) The SDs of FDP, FNP, and TP for the cube with about $10\%$ signal proportion.
(c) The means of FDP, FNP, and TP for the cube with about $20\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.51.
(d) The SDs of FDP, FNP, and TP for the cube with about $20\%$ signal proportion.
(e) The means of FDP, FNP, and TP for the cube with about $30\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.45.
(f) The SDs of FDP, FNP, and TP for the cube with about $30\%$ signal proportion.
Figure 3: Simulation results over 50 replications at a nominal FDR level of 0.05.
(a) The means of FDP, FNP, and TP for the cube with about $10\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.58.
(b) The SDs of FDP, FNP, and TP for the cube with about $10\%$ signal proportion.
(c) The means of FDP, FNP, and TP for the cube with about $20\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.42.
(d) The SDs of FDP, FNP, and TP for the cube with about $20\%$ signal proportion.
(e) The means of FDP, FNP, and TP for the cube with about $30\%$ signal proportion. The mean FDP values of NeuralFDR are omitted because they exceed 0.44.
(f) The SDs of FDP, FNP, and TP for the cube with about $30\%$ signal proportion.
Figure 4: Simulation results over 50 replications at a nominal FDR level of 0.1.
3.2 Simulation Settings
We generate each simulated dataset on a $30\times 30\times 30$ lattice cube, yielding $m=27{,}000$ voxels. Since it is computationally prohibitive to generate the true hypothesis states $\boldsymbol{h}=(h_{1},\dots,h_{m})$ from the fully connected Markov random field in (5) using the Gibbs sampler, which requires $O(m^{2}B)$ time for $B$ burn-in iterations, we follow [21] and derive $\boldsymbol{h}$ from the ADNI FDG-PET dataset described in Section 4.2. Specifically, we use the result of the q-value method at a nominal FDR level of 0.01 for the comparison between the cognitively normal group and the group of patients with late mild cognitive impairment who later converted to AD. From the brain volume of the q-value result, we randomly crop three $30\times 30\times 30$ lattice cubes, each containing approximately 10%, 20%, or 30% of voxels identified as significant. In each cube, we assign $h_{i}=1$ to significant voxels and $h_{i}=0$ otherwise. To enhance spatial coherence, isolated signals are reset to $h_{i}=0$, leaving the signal proportion within each cube nearly unchanged. The test statistics $\boldsymbol{x}=(x_{1},\dots,x_{m})$ are generated from the Gaussian mixture model: $x_{i}|h_{i}\sim(1-h_{i})\mathcal{N}(0,1)+h_{i}\{0.5\mathcal{N}(\mu_{1},\sigma_{1}^{2})+0.5\mathcal{N}(2,1)\}$. We consider two types of settings: (i) varying $\mu_{1}$ from $-4$ to 0 with $\sigma_{1}^{2}=1$, and (ii) varying $\sigma_{1}^{2}$ from 0.125 to 8 with $\mu_{1}=-2$. The mean difference $\Delta\mu_{i}$ used for the appearance kernel in (6) is estimated from the ADNI FDG-PET data as the sample mean difference adjusted for covariates via linear regression (see Section 4.2). In total, we consider 45 simulation settings (15 per signal proportion level). We conduct the nine FDR control methods at a nominal FDR level $\alpha\in\{0.05,0.1\}$ for 50 independent replications of each simulation setting.
3.3 Evaluation Metrics
We evaluate the performance of the nine FDR control methods in multiple testing by reporting the means and SDs of the FDP, FNP, and TP over the 50 simulation replications, where the mean FDP and mean FNP are empirical estimates of the FDR and FNR, respectively. Since valid FDR control is a fundamental requirement for FDR control methods, ensuring that the mean FDP over replications (i.e., the empirical FDR) is well controlled serves as a prerequisite for considering the other metrics. In addition, we assess the timing performance of each method by reporting the means and SDs of its runtime.
3.4 Multiple Testing Results
The multiple testing results are summarized in Figures 3 and 4. As shown in subfigures (a), (c) and (e) of both figures, our proposed fcHMRF-LIS successfully controls the empirical FDR (i.e., the mean FDP) at the nominal level in all simulation settings. LAWS and LocalFDR also achieve valid empirical FDR control in all settings, and BH and q-value do so in nearly all settings. In contrast, NeuralFDR and nnHMRF-LIS fail to control the empirical FDR in all settings, SmoothFDR fails in multiple settings, and DeepFDR fails in several. Moreover, fcHMRF-LIS achieves the highest mean TP in 42 out of 45 settings at both nominal FDR levels (0.05 and 0.1), among methods that successfully control the empirical FDR in each setting. The three exceptions occur in the cube with approximately 20% signal proportion and variance $\sigma_{1}^{2}=1$: for nominal FDR level 0.05 at $\mu_{1}\in\{-1.5,-1,-0.5\}$ (Figure 3(c)), and for level 0.1 at $\mu_{1}\in\{-1,-0.5,0\}$ (Figure 4(c)). In these settings, fcHMRF-LIS is outperformed by DeepFDR but still ranks second in mean TP among the methods with empirical FDR under control. In terms of empirical FNR (i.e., the mean FNP), and still comparing only among methods that successfully control the empirical FDR per setting, fcHMRF-LIS ranks third (only behind DeepFDR and LAWS) in the three exceptions at each nominal FDR level, ranks first in 11 settings at level 0.05 and in 36 settings at level 0.1, and ranks second in the remaining settings. Theoretically, if the data exactly follow our fcHMRF model, the LIS-based testing procedure (LABEL:def:_LIS_procedure) asymptotically achieves the smallest FNR among all valid FDR control methods [12]. Since the simulated data are based on the real ADNI data, they may not perfectly satisfy the fcHMRF model, which may lead to suboptimal performance of fcHMRF-LIS in some settings.
However, a key strength of fcHMRF-LIS lies in its stability across replications. In contrast, DeepFDR and LAWS exhibit substantial variability, particularly in the aforementioned settings where fcHMRF-LIS shows suboptimality in mean TP and mean FNP. As shown in subfigures (b), (d) and (f) of Figures 3 and 4, all four recent spatial methods (DeepFDR, LAWS, SmoothFDR, and NeuralFDR) display high variability in FDP, FNP, and/or TP across all or some settings. In comparison, fcHMRF-LIS consistently achieves low SDs of FDP, FNP, and TP in all settings. The nearest-neighbor-based method, nnHMRF-LIS, also demonstrates strong stability in FNP and TP, though it fails to control the empirical FDR. The classical non-spatial methods BH, q-value, and LocalFDR prioritize FDR control but have substantially lower statistical power compared to the spatial methods fcHMRF-LIS, DeepFDR, and LAWS. Nevertheless, these three classical methods exhibit strong stability, with low variability in FNP and TP across all settings and in FDP in most settings, comparable to that of our fcHMRF-LIS.
3.5 Timing Performance
We evaluate the timing performance of the nine FDR control methods. BH, q-value, LocalFDR, SmoothFDR, LAWS, and fcHMRF-LIS are run on a server with 20 Intel Xeon Platinum 8268 CPU cores (2.90GHz, total 64GB RAM). DeepFDR, NeuralFDR, and nnHMRF-LIS are executed on the same CPU configuration and an NVIDIA RTX8000 GPU (48GB VRAM). Runtime is measured under the simulation setting with $(\mu_{1},\sigma_{1}^{2})=(-2,1)$ and approximately $20\%$ signal proportion at the nominal FDR level of 0.1. Table 2 reports the mean and SD of runtime over the 50 simulation replications for each method. As expected, BH, q-value, and LocalFDR, which are designed for independent tests rather than spatial data, achieve the fastest performance, each completing in less than 0.14 seconds on average with minimal variability. Our fcHMRF-LIS ranks seventh in speed, with a mean runtime of 221.14 seconds and an SD of 21.83 seconds. This is approximately 5/8 the runtime of LAWS and about 1/27 that of NeuralFDR. In comparison, fcHMRF-LIS takes about 2.4 times the runtime of the simpler nnHMRF-LIS method and about 1.7 times that of SmoothFDR. However, as the number of tests increases from 27,000 in the simulation setting to 439,758 in the real ADNI dataset, fcHMRF-LIS exhibits significantly better scalability than nnHMRF-LIS and SmoothFDR, requiring only about 1/5 of the runtime of nnHMRF-LIS and about 1/14 that of SmoothFDR; see Section 4.4 for further details.
Table 2: Mean (SD) of runtime in seconds.
Method Simulation ADNI
BH 0.050 (0.001) 0.080 (0.015)
q-value 0.135 (0.003) 2.203 (0.043)
LocalFDR 0.036 (0.005) 0.260 (0.036)
nnHMRF-LIS 92.19 (10.23) 20245 (1987)
SmoothFDR 133.22 (8.16) 58745 (7397)
NeuralFDR 5995 (442.4) 92589 (13361)
LAWS 349.38 (29.19) 538146 (15102)
DeepFDR 7.708 (1.617) 89.98 (5.17)
fcHMRF-LIS 221.14 (21.83) 4341 (502.4)
Figure 5: The z-statistics of discoveries for EMCI2AD vs. CN.
Figure 6: The z-statistics of discoveries for LMCI2AD vs. CN.
Figure 7: The z-statistics of discoveries for AD vs. CN.
(a) EMCI2AD vs. CN
(b) LMCI2AD vs. CN
(c) AD vs. CN
Figure 8: Heatmaps illustrating the discovery proportion within each ROI for the three comparisons in the ADNI data analysis.
Figure 9: Top 20 ROIs identified by fcHMRF-LIS and discovery proportions from nine methods for each of the three comparisons in the ADNI data analysis.
4 ADNI FDG-PET Data Analysis
4.1 Study Aim
AD is the most common form of dementia, affecting more than 30 million people worldwide, and there is still no curative treatment [46]. FDG-PET, which measures brain glucose metabolism, is widely used to support the clinical diagnosis of AD and to monitor its progression [5]. Because glucose metabolism in the brain typically declines with AD, differences across disease stages can be examined by testing voxel-wise mean differences in SUVR. This yields a high-dimensional spatial multiple-testing problem. In this study, we apply voxel-wise multiple testing to compare the mean SUVR difference between the cognitively normal (CN) group and each of the following three groups: early mild cognitive impairment (MCI) patients with conversion to AD (EMCI2AD), late MCI patients with conversion to AD (LMCI2AD), and the AD group. We evaluate and compare the nine FDR control methods given in Section 3.1.
4.2 ADNI FDG-PET Data
We use the FDG-PET brain image dataset obtained from the ADNI database (https://adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging (MRI), positron emission tomography (PET), other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of MCI and early AD. We use baseline 3D FDG-PET scans from 742 subjects, including 286 CN subjects, 42 EMCI2AD patients, 175 LMCI2AD patients, and 239 AD patients. All scans are preprocessed using the Clinica software [47], which includes spatial normalization to the MNI IXI549Space template and intensity normalization based on average uptake in the pons region. We consider the 120 regions of interest (ROIs) defined by the AAL2 atlas [48], resulting in a total of $m=439{,}758$ voxels. At each voxel, we fit a linear regression model with SUVR as the response, and dummy variables for the EMCI2AD, LMCI2AD, and AD groups as predictors, using the CN group as the reference. The model adjusts for age, gender, race, ethnicity, education, marital status, and APOE4 status by including them as covariates. From this model, we obtain the voxel-level t-statistics and p-values for the regression coefficients $\{\hat{\beta}_{\text{EMCI2AD},i},\hat{\beta}_{\text{LMCI2AD},i},\hat{\beta}_{\text{AD},i}\}_{i=1}^{m}$, corresponding to the dummy variables for the three AD-related stage groups. These coefficients represent the three comparisons: EMCI2AD vs. CN, LMCI2AD vs. CN, and AD vs. CN. The voxel-level mean difference between an AD-related stage and CN is estimated by $\Delta\hat{\mu}_{i}=\hat{\beta}_{\text{stage},i}$. All t-statistics are converted to z-statistics for use in FDR control methods that require them as input.
4.3 Multiple Testing Results
Nominal FDR levels of $\alpha=0.05$, 0.01, and 0.005 are applied to the EMCI2AD vs. CN, LMCI2AD vs. CN, and AD vs. CN comparisons, respectively. Stricter nominal levels are used for comparisons involving later disease stages to better control the number of false discoveries. Table 3 summarizes the number of discoveries made by each method and the proportions of discoveries with p-values exceeding the nominal level and 0.05. Figures 57 present the z-statistics of discoveries for each method. Most detected brain areas exhibit hypometabolism, with affected areas expanding as AD progresses. This trend is further illustrated in Figure 8, which shows heatmaps of the discovery proportion in each ROI for all methods.
The EMCI2AD vs. CN comparison is the most challenging among the three comparisons, but is also the most critical for enabling early detection and timely therapeutic intervention in AD. In this comparison, BH, q-value, LocalFDR, and LAWS yield no discoveries. In fact, the three classical non-spatial methods BH, q-value, and LocalFDR fail to produce any discoveries until the nominal FDR level is raised to 0.8. As shown in Table 3, nnHMRF-LIS, SmoothFDR, NeuralFDR, DeepFDR, and fcHMRF-LIS identify 13587, 141656, 25675, 11341, and 20758 discoveries, respectively. However, a substantial proportion of discoveries made by nnHMRF-LIS (55.86%), SmoothFDR (80.48%), and NeuralFDR (95.10%) have p-values greater than the nominal level of 0.05. While a discovery with a p-value above the nominal level is not necessarily a false signal, a large proportion of such discoveries can raise concerns about the validity of these methods in controlling the FDR. In contrast, DeepFDR and fcHMRF-LIS report only 2.14% and 4.75% of discoveries with p-values greater than 0.05, respectively, suggesting more credible results. Among all methods except SmoothFDR and NeuralFDR, fcHMRF-LIS achieves the largest number of discoveries.
In the LMCI2AD vs. CN comparison, fcHMRF-LIS identifies 181209 discoveries, with 13.38% having p-values greater than the nominal level of 0.01 and 0.82% exceeding 0.05. Again, it produces the most discoveries aside from SmoothFDR and NeuralFDR, both of which yield more than 42% of discoveries with p-values greater than 0.01 and over 27% exceeding 0.05. The distribution of discoveries from fcHMRF-LIS is similar to those from the six other methods excluding SmoothFDR and NeuralFDR, as shown in Figures 6 and 8(b).
In the AD vs. CN comparison, fcHMRF-LIS detects 260229 discoveries, with 3.36% having p-values greater than the nominal level of 0.005 and 0.04% exceeding 0.05. Its distribution of discoveries is again similar to those obtained by the six methods other than SmoothFDR and NeuralFDR, as shown in Figures 7 and 8(c). In contrast, SmoothFDR and NeuralFDR report 24.40% and 37.54% of discoveries with p-values greater than 0.005, and 13.35% and 25.21% exceeding 0.05, respectively.
Across all three comparisons, SmoothFDR and NeuralFDR appear to overestimate the signals, as evidenced by the large proportions of discoveries with p-values exceeding the nominal levels and even 0.05. Additionally, Figures 57 reveal that SmoothFDR tends to produce oversmoothed results, while NeuralFDR yields discoveries with reduced spatial coherence.
Figure 9 shows the top 20 ROIs for each of the three comparisons, ranked by discovery proportions from fcHMRF-LIS, along with corresponding values from all nine methods. For the EMCI2AD vs. CN comparison, discovery proportions vary widely across the nine methods. This variability decreases substantially for the LMCI2AD vs. CN comparison and becomes even smaller for the AD vs. CN comparison, indicating increasing agreement among methods on the most affected ROIs in later disease stages. In the most challenging EMCI2AD vs. CN comparison, as noted earlier, our fcHMRF-LIS yields the largest number of high-quality discoveries. Its top 20 ROIs include lobules VII, VIII, and IX of the vermis, bilateral parahippocampal gyri, bilateral hippocampi, bilateral paracentral lobules, bilateral precentral gyri, bilateral superior temporal poles, left middle temporal pole, left middle temporal gyrus, left inferior temporal gyrus, left amygdala, left fusiform gyrus, left postcentral gyrus, and left lobule IX of the cerebellar hemisphere. These 20 ROIs are not detected in the FDG-PET data by any of the three classical methods (BH, q-value, and LocalFDR) or by the recent spatial method LAWS. However, prior studies using structural or functional MRI data have shown that these regions are affected in early stages of AD [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], providing strong support for the neurobiological relevance and enhanced sensitivity of the fcHMRF-LIS method.
4.4 Time and Memory Efficiency
We execute the nine FDR control methods on the same computational resource as specified in Section 3.5. Table 2 reports each method’s mean and SD of runtime over the three comparisons in the ADNI data analysis. Our fcHMRF-LIS achieves a mean runtime of 4341 seconds (SD = 502.4 seconds), ranking fifth overall, only behind the three non-spatial methods (BH, q-value, and LocalFDR) and the deep learning-based method DeepFDR. As expected, the three non-spatial methods are the fastest, each with a mean runtime under 2.3 seconds (SD under 0.05 seconds), as they do not account for spatial dependence among tests. DeepFDR benefits from GPU acceleration and, like most deep learning models, is expected to exhibit substantially longer runtime when executed without GPU support. In contrast, our fcHMRF-LIS runs entirely on CPUs and is significantly more efficient than other spatial methods, requiring only about 1/5 the runtime of nnHMRF-LIS, 1/14 of SmoothFDR, 1/21 of NeuralFDR, and 1/124 of LAWS.
Table 4 presents the mean and SD of peak memory usage for each method over the three comparisons. The non-spatial methods BH, q-value, and LocalFDR exhibit minimal CPU memory usage (mean $\leq$ 0.56 GB, SD $\leq$ 0.07 GB) and require no GPU. Between the two deep learning-based methods, DeepFDR shows the highest GPU memory usage (mean = 33.75 GB, SD = 0.91 GB) due to its W-Net architecture, whereas NeuralFDR, based on a multilayer perceptron, consumes less GPU memory (mean = 15.39 GB, SD = 0.81 GB) but requires substantially longer runtime. Our fcHMRF-LIS demonstrates moderate peak CPU memory usage (mean = 2.91 GB, SD = 0.59 GB) while operating entirely on CPUs. All methods can be executed on standard hardware, with minimal memory requirements for the non-spatial methods and moderate demands for the spatial methods. Overall, our fcHMRF-LIS achieves excellent computational efficiency and scalability for high-dimensional data analysis.
Table 3: Summary of discoveries for the three comparisons in ADNI data analysis. Abbreviations: Disc. = discoveries; w/ = with; $p$ = p-value.
Method #$\#\{\text{Disc.}\}$. } #{Disc. w/ $\dfrac{\#\{\text{Disc. w/ }p{>}\alpha\}}{\#\{\text{Disc.}\}}${Disc. w/ $\dfrac{\#\{\text{Disc. w/ }p{>}0.05\}}{\#\{\text{Disc.}\}}$
EMCI2AD vs. CN ($\alpha=0.05$; #$\#\{p\leq\alpha\}=27645$)
BH 0
q-value 0
LocalFDR 0
nnHMRF-LIS 13587 0.5586 0.5586
SmoothFDR 141656 0.8048 0.8048
NeuralFDR 25675 0.9510 0.9510
LAWS 0
DeepFDR 11341 0.0214 0.0214
fcHMRF-LIS 20758 0.0475 0.0475
LMCI2AD vs. CN ($\alpha=0.01$; #$\#\{p\leq\alpha\}=157075$)
BH 119374 0.0000 0.0000
q-value 159516 0.0153 0.0000
LocalFDR 117783 0.0000 0.0000
nnHMRF-LIS 155592 0.0400 0.0093
SmoothFDR 256271 0.4240 0.2797
NeuralFDR 255743 0.5375 0.4232
LAWS 125562 0.0007 0.0000
DeepFDR 115325 0.0034 0.0000
fcHMRF-LIS 181209 0.1338 0.0082
AD vs. CN ($\alpha=0.005$; #$\#\{p\leq\alpha\}=251543$)
BH 240571 0.0000 0.0000
q-value 282201 0.1086 0.0000
LocalFDR 290968 0.1390 0.0000
nnHMRF-LIS 252794 0.0227 0.0026
SmoothFDR 326667 0.2440 0.1335
NeuralFDR 375010 0.3754 0.2521
LAWS 236787 0.0042 0.0000
DeepFDR 227756 0.0063 0.0002
fcHMRF-LIS 260229 0.0336 0.0004
Table 4: Mean (SD) of peak memory usage (in GB) for the ADNI data analysis.
Method RAM (CPU) VRAM (GPU)
BH 0.469 (0.052)
q-value 0.471 (0.056)
LocalFDR 0.557 (0.062)
nnHMRF-LIS 0.911 (0.085) 0.302 (0.022)
SmoothFDR 3.380 (0.401)
NeuralFDR 0.766 (0.068) 15.385 (0.814)
LAWS 2.413 (0.349)
DeepFDR 3.026 (0.570) 33.752 (0.908)
fcHMRF-LIS 2.909 (0.592)
5 Conclusion
In this paper, we propose fcHMRF-LIS, a novel spatial FDR control method for large-scale voxel-wise multiple testing in neuroimaging studies. The method is designed to address three major limitations of existing methods: inadequate modeling of spatial dependence, high variability in FDP and FNP, and limited computational scalability. To overcome these challenges, fcHMRF-LIS integrates the LIS-based testing procedure with our newly designed fcHMRF, enabling effective and efficient inference under complex spatial structures.
We develop an efficient EM algorithm for parameter estimation using the mean-field approximation, CRF-RNN technique, and permutohedral lattice filtering, reducing time complexity from quadratic to linear in the number of tests. In extensive simulations, fcHMRF-LIS achieves superior FDR control, lower FNR, higher TP, and reduced variability across replications, compared to eight existing methods, including classical, local smoothing-based, and deep learning-based approaches.
When applied to ADNI FDG-PET data containing 439,758 voxels of interest, fcHMRF-LIS completes each of three comparisons in about 1.2 hours on a 20-core CPU server without GPU acceleration. This is significantly faster than LAWS, which requires about 149.5 hours on the same server, and more efficient than nnHMRF-LIS and NeuralFDR, which take about 5.6 and 25.7 hours, respectively, using a GPU. Notably, in the most challenging EMCI2AD vs. CN comparison, fcHMRF-LIS identifies 20 ROIs that are supported by prior structural and functional MRI studies, but missed by classical FDR control methods using the FDG-PET data. This highlights both the sensitivity and neurobiological relevance of fcHMRF-LIS.
These results demonstrate that fcHMRF-LIS is a powerful, stable, and scalable method for voxel-wise multiple testing in neuroimaging. Future work will explore extensions of this framework to other domains with complex spatial or sequential structures, such as GWAS SNP data and time series data.
Acknowledgements
Dr. Shu’s research was partially supported by NYU GPH Goddard Award and NYU GPH Research Support Grant. Dr. de Leon’s research was partially supported by the grants R01 AG12101, R01 AG022374, R01 AG13616, RF1 AG057570, and R56 AG058913 from the National Institutes of Health (NIH). The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
Data collection and sharing for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904) and DOD ADNI (Department of Defense award number W81XWH-12-2-0012). ADNI is funded by the National Institute on Aging, the National Institute of Biomedical Imaging and Bioengineering, and through generous contributions from the following: AbbVie, Alzheimer’s Association; Alzheimer’s Drug Discovery Foundation; Araclon Biotech; BioClinica, Inc.; Biogen; Bristol-Myers Squibb Company; CereSpir, Inc.; Cogstate; Eisai Inc.; Elan Pharmaceuticals, Inc.; Eli Lilly and Company; EuroImmun; F. Hoffmann-La Roche Ltd and its affiliated company Genentech, Inc.; Fujirebio; GE Healthcare; IXICO Ltd.; Janssen Alzheimer Immunotherapy Research & Development, LLC.; Johnson & Johnson Pharmaceutical Research & Development LLC.; Lumosity; Lundbeck; Merck & Co., Inc.; Meso Scale Diagnostics, LLC.; NeuroRx Research; Neurotrack Technologies; Novartis Pharmaceuticals Corporation; Pfizer Inc.; Piramal Imaging; Servier; Takeda Pharmaceutical Company; and Transition Therapeutics. The Canadian Institutes of Health Research is providing funds to support ADNI clinical sites in Canada. Private sector contributions are facilitated by the Foundation for the National Institutes of Health (www.fnih.org). The grantee organization is the Northern California Institute for Research and Education, and the study is coordinated by the Alzheimer’s Therapeutic Research Institute at the University of Southern California. ADNI data are disseminated by the Laboratory for Neuro Imaging at the University of Southern California.
Appendix A Implementation Details of Methods
This appendix provides implementation details for all methods used in our simulation studies and ADNI data analysis. Python implementations were used whenever available; otherwise, R versions were employed. All numerical studies were conducted using Python 3.12.5 and R 4.2.1.
  • BH: Takes a 1D sequence of p-values as input. We used the Python package statsmodels (v0.12.2), available at https://www.statsmodels.org.
  • q-value: Takes a 1D sequence of p-values as input. We used the Python package multipy (v0.16), available at https://github.com/puolival/multipy.
  • LocalFDR: Takes a 1D sequence of z-statistics as input. Implemented using the Python package statsmodels (v0.12.2), available at https://www.statsmodels.org.
  • nnHMRF-LIS: Take a 3D volume of z-statistics as input. We used the GPU-enabled Python package from https://github.com/shu-hai/nnHMRF-LIS. In simulations, a single HMRF was used for the $30\times 30\times 30$ lattice cube. In ADNI data analysis, each ROI was modeled with a separate HMRF, and the LIS values were then pooled for testing.
  • SmoothFDR: Take a 3D volume of z-statistics as input. We used the Python package from https://github.com/tansey/smoothfdr, with 20 sweeps.
  • NeuralFDR: Takes p-values as input along with 3D spatial coordinates as covariates. Implemented using the Python package from https://github.com/fxia22/NeuralFDR. For simulations, default parameters were used. For ADNI data, we modified the code to enable mini-batch training (batch size=27,000) due to GPU memory constraints, and set the parameters n-init=3 and num-iterations=200.
  • LAWS: Takes a 3D volume of p-values as input. We used the R code provided in the supplementary materials of the LAWS paper [18], available at https://doi.org/10.1080/01621459.2020.1859379.
  • DeepFDR: Takes a 3D volume of test statistics (z-statistics for simulations and t-statistics for ADNI data) and corresponding 3D volume of p-values as input. We used the Python package from https://github.com/kimtae55/DeepFDR. For simulations, the $30\times 30\times 30$ volume was zero-padded to $32\times 32\times 32$ to facilitate the two max-pooling layers in each U-net of the DeepFDR network.
  • fcHMRF-LIS: Takes a 3D volume of z-statistics and corresponding 3D volume of mean-difference estimates $\{\Delta\hat{\mu}_{i}\}_{i=1}^{m}$ as input. The Python package implementing the fcHMRF-LIS method is available at https://github.com/kimtae55/fcHMRF-LIS. The parameter vector $\boldsymbol{w}=(w_{0},w_{1},w_{2})$ was initialized as $(0.5,1,1)$ by default. In simulation settings with weak signals $(\mu_{1}\in\{0,-0.5,-1\})$, we found that initializing $w_{2}=5$ improved the spatial coherence of discoveries. Thus, we instead initialized $w_{2}=5$ in these cases. The non-null distribution $f_{1}$ was initialized using (11) with $q_{i}(h_{i}=1|\boldsymbol{x})$ set to $1-$p-valuei. For ADNI data, to reduce computational burden, 50% of voxels were randomly sampled to compute $\operatorname{SD}(\{\delta_{ij}\}_{i\neq j})$ where $\delta_{ij}\in\{\ell_{i,s}-\ell_{j,s},\Delta\mu_{i}-\Delta\mu_{j}\}$; The resulting standard deviations were used to define the corresponding kernel bandwidth parameters $\{\theta_{\alpha,s},\theta_{\gamma,s},\theta_{\beta}\}$. The number of EM iterations was set to 25, with early stopping activated based on the loss $-\hat{Q}(\boldsymbol{\phi}^{(t+1)}|\boldsymbol{\phi}^{(t)})$ and a patience threshold of 5. In each EM iteration, we minimized $-\hat{Q}_{2}$ in (2.4) using the AdamW optimizer with an initial learning rate of $10^{-4}$ and a weight decay of 0.01, running for five epochs. The mean-field approximation in (14)-(16) was performed with $R=5$ iterations each time it was applied. To compute $\hat{Q}_{2}$ in (2.4), we generated a moderate number of samples $\{h_{i}^{(n,t)}\}_{n=1}^{N=100}$.
In the ADNI data analysis, the original brain image volume of size $121\times 145\times 121$ was cropped to $100\times 120\times 100$ to tightly enclose the ROIs. Non-ROI voxels were set to 0 for t-statistics and z-statistics, and to 1 for p-values. Multiple testing was performed only on the ROI voxels, ensuring that non-ROI voxels were excluded from the analysis.